OpenStudy (nathan917):

From a committee of 6 girls and 4 boys, a name is selected. Then another name is selected. What is the probability that both people drawn will be girls? nine over twenty-five two-fifteenths three-fifths one-third

OpenStudy (rulnick):

Prob both girls = Prob first is girl x Prob second is girl given first was girl.

OpenStudy (nathan917):

probability of drawing a girl in the first pick x probability of drawing a girl in the 2nd pick = 6/10 x 5/9 = (6x5) / (10x9) = 30 / 90 = 1 / 3

OpenStudy (rulnick):

Yes!

OpenStudy (nathan917):

ok. Thanks.

OpenStudy (rulnick):

welcome

OpenStudy (goldphenoix):